a traffic light is red for 25 sec, yellow for 5 sec, and green for 55 sec. What is the probability that when you reach the light, the light is yellow
Answer by stanbon(75887)

a traffic light is red for 25 sec, yellow for 5 sec, and green for 55 sec. What is the probability that when you reach the light, the light is yellow
---
yellow time: 5 sec
total recycle time: 85 sec
---
P(yellow) = 5/85 = 1/17
==============================
Cheers,
Stan H.